OpenStudy (anonymous):

What is the area of a square with side lengths of 7 feet? A. 28 ft B. 28 ft2 C. 49 ft2 D. 49 ft

OpenStudy (anonymous):

49 sq feet

OpenStudy (anonymous):

so 49ft^2

OpenStudy (yttrium):

Just be reminded of this formula. A = s^2 A = s*s A = 7*7 A=?

OpenStudy (yttrium):

@samsterz , don't just answer it. Be reminded of our code of conduct.

OpenStudy (anonymous):

its... C

OpenStudy (anonymous):

7 feet^2 B times h = A

OpenStudy (anonymous):

7 feet^2 = 49 feet^2

OpenStudy (anonymous):

yup as i said its c

OpenStudy (yttrium):

People please read http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i explained it yttrium this is pretty self explan... work

OpenStudy (yttrium):

but you answered it first before explaining. Just be careful next time